OpenStudy (unklerhaukus):

Given 100 multiple choice questions, what is the probability of getting 50 questions right, if the selection A, B, C, or D is random.

OpenStudy (amistre64):

\[\binom{n}{k}p^kq^{n-k}\]

ganeshie8 (ganeshie8):

binomial distribution

OpenStudy (anonymous):

Multiple*

OpenStudy (anonymous):

That equation provided by @amistre64 should pretty much help you solve it.

OpenStudy (unklerhaukus):

\[\binom{100}{50}(\tfrac14)^{100}q^{100-50}\] What is q?

OpenStudy (amistre64):

q is the complement of p ...

OpenStudy (amistre64):

if we have 1/4 chance of success, whats the alternative?

OpenStudy (amistre64):

(p+q)^n; such that p+q = 1 for a valid distribution. therefore q = 1-p

OpenStudy (unklerhaukus):

3/4

OpenStudy (amistre64):

we can expand that to (p+q+r+...+z)^n such that p+q+r+...+z=1 and we can determine a lot more variables :) yes, q = 3/4

OpenStudy (unklerhaukus):

\[\binom{100}{50}(\tfrac14)^{100}(\tfrac14-1)^{100-50}\\~\\ =^{100}C_{50}\times\tfrac1{4^{100}}\times(1-\tfrac14)^{100-50}\\ =\frac{100!}{50!(100-50)!}\times4^{-100}\times(\tfrac34)^{50}\\ =\frac{100!}{(50!)^2}\times4^{-150}\times3^{50}\\ \approx 3.6\times10^{-38}\]

OpenStudy (unklerhaukus):

~ so it is improbable , thank you

OpenStudy (unklerhaukus):

is this for getting exactly 50 or at least 50?

ganeshie8 (ganeshie8):

exactly 50

OpenStudy (unklerhaukus):

*\[\binom{100}{50}(\tfrac14)^{50}(\tfrac14-1)^{100-50}\\~\qquad\qquad\quad\vdots\\ =\frac{100!}{(50!)^2}\times4^{-100}\times3^{50}\\\]

ganeshie8 (ganeshie8):

atleast 50 is also very improbable

ganeshie8 (ganeshie8):

atleast 25 gives almost half the area :

OpenStudy (unklerhaukus):

i see, but how do i modify the equation \[\sum_{k=50}^{100}\binom{100}{k}(\tfrac14)^{k}(\tfrac14-1)^{k-50}\] ?

ganeshie8 (ganeshie8):

more than half *

OpenStudy (unklerhaukus):

yeah i would have expect to get 25 right

ganeshie8 (ganeshie8):

we may play with combinatorics, but not sure if we can simplify it.... we use normal approximation when n is large (n > 30)

ganeshie8 (ganeshie8):

we convert the binomial distribution into normal distribution and use the zscore tables to extract the areas

ganeshie8 (ganeshie8):

For our present problem, normal approximation would be : mean \(\mu = np = 100*1/4 = 25 \) standard deviation \(\sigma = \sqrt{np(1-p)} = \sqrt{100*1/4*3/4} = 5\sqrt{3}/2\)

ganeshie8 (ganeshie8):

Zscore for observation 50 = \(\large \dfrac{50 -26 }{5\sqrt{3}/2} = 5.54 \)

ganeshie8 (ganeshie8):

P(X>= 50) : http://www.wolframalpha.com/input/?i=P%28X+%3E%3D++5.54%29 this is the normal approximation forearlier summation formula : \(\sum_{k=50}^{100}\binom{100}{k}(\tfrac14)^{k}(\tfrac14-1)^{k-50}\)

ganeshie8 (ganeshie8):

It really matches quite closely as n is large (100)

ganeshie8 (ganeshie8):

the larger the value of n, the lesser is the error...